Grade 12th passPhysical Chemistry

Calculate the volume occupied by 16kg of chlorine at pressure 743 mm Hg and 20 cectegrade

Surajit Jana
7 Years agoGrade 12th pass

1 Answer

Pooja
7 Years ago
Pressure is 743 mmHg
P= 743÷760 atm
Mass is 16 kg
No of moles(n)=(16×1000)÷71
Temperature=20℃
T=20+273=293K
Now PV=nRT
V=nRT/P
    ={(16×1000÷71) × 0.0821 ×293}÷743/760
  =5.54 kiloliters